• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 10:49:14 on October 18, 2010

    Entry number 335453

    keyword=Status update

    Currently - all halls are down for work on the LCW system. Estimated restoration of water flow is about 7 pm. In the meantime, we will remain in Controlled Access.

    Work on the Hall C Moller cryo re-sting is in progress - impact on the target is already evident.

    When we are ready for beam this evening - the present plan is to continue with elastic calibration data.